J. Y Limited consulted Mr Bansah to estimate their supply function. Mr Bansah identified their intercept as - 4 and also indicated that the firms supply increases by 2units each time price increases by a cedi. Write down the equation for J. Y's supply function.


1
Expert's answer
2021-02-22T17:03:39-0500

 P=a + bS

a = plots the starting point of the supply curve on the Y-axis intercept.

b = slope of the supply curve

Slope=1/2=0.5

P=-4+0.5Qs